# EUROPEAN ORGANIZATION FOR NUCLEAR RESEARCH (CERN)

### v1

@@ 4 Fit to the dipion spectrum
There is a long history of possible vector states in the high-mass region as discussed in the PDG review [15]. The $\rho^{\prime}$ resonance is measured to have a mass of $1350\pm 5\pm 20\text{\,Me\kern-1.00006ptV}$ and a width of $320\pm 10\pm 40\text{\,Me\kern-1.00006ptV}$, while the $\rho^{\prime\prime}$ resonance has a mass of $1790\pm 5\pm 20\text{\,Me\kern-1.00006ptV}$ and a width of $290\pm 10\pm 40\text{\,Me\kern-1.00006ptV}$, where the first uncertainty is the statistical precision of the fit result and the second is due to systematic effects, evaluated using the standard deviation of results obtained by fitting the model separately in each rapidity bin. However, it should be pointed out that improvements to the fit have been obtained with alternative parametrisations including using a form factor to multiply $\cal{BW}_{\rho}$, and including a contribution from a noninterfering $f_{2}(1270)$ meson.
These measurements are consistent with the values of the $\rho(1450)$ and $\rho(1700)$ mesons in the PDG [15], noting the role the choice of dynamic lineshapes can play. Further evidence for the existence of these particles was reported by the Babar collaboration using $e^{+}e^{-}\rightarrow\pi\pi(\gamma)$ data [48] where the same qualitative features are observed as in the LHCb data. The production processes are similar with the main difference being that in photoproduction the photon is quasi-real while it is strongly virtual in $e^{+}e^{-}$ annihilation. The coupling between the photon and the pions is probed in both processes so the LHCb data may help clarify the resonant contributions, which have an impact in determining the hadronic corrections to the $g-2$ measurement [49]. In Ref. [25], the ALICE collaboration reports an excess of events in the high-mass region, corresponding to a resonance mass of $1725\pm 17\text{\,Me\kern-1.00006ptV}$ and a width of $143\pm 21\text{\,Me\kern-1.00006ptV}$, as determined with a simple Gaussian model. With a similar model, the STAR collaboration [47] obtained a mass of $1653\pm 10\text{\,Me\kern-1.00006ptV}$ and a width of $164\pm 15\text{\,Me\kern-1.00006ptV}$, which they noted was compatible with the spin-3 $\rho_{3}(1690)$.
@@ 5 Cross-section
where $N_{V}$ is the number of vector mesons associated to the signal component $\cal{S}$, the integrated luminosity is represented by ${\cal L}$, the bin width by $\Delta_{y}$ and ${\rm BF}_{V\rightarrow\pi\pi}$ is the branching fraction of the decay of vector meson, $V$, to two charged pions. For the $\rho$ meson, ${\rm BF}_{\rho\rightarrow\pi\pi}$ is taken to be $0.9901\pm 0.0016$ by subtracting the branching fractions for the sub-dominant decay modes [15] from unity, while for the $\omega$ meson ${\rm BF}_{\omega\rightarrow\pi\pi}=0.0153\pm 0.0012$. For the $\rho^{\prime}$ and $\rho^{\prime\prime}$ mesons, where the branching fractions are not measured, the product ${\rm BF}_{V\rightarrow\pi\pi}(\mathrm{d}\sigma_{V}/{\mathrm{d}y})$ is reported. A tiny correction, $f_{\rm smog}=0.9985\pm 0.0001$ is applied to account for vector meson photoproduction in fixed-target collisions caused by the injection of gas into the interaction region needed for a precision luminosity measurement [36].
@@
Table 5: Differential cross-section for the $\rho,\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$ mesons in bins of rapidity. The first uncertainty includes statistical and uncorrelated systematics. The second uncertainty represents the correlated systematics.
@@
Figure 5: Differential cross-section for coherent $\rho$ photoproduction as a function of rapidity. The LHCb error bars represent the statistical and total uncertainties. The shaded band corresponds to the estimated uncertainty on the GKZ prediction. The contributions due to the photon emitted from each ion are labelled as $W^{+}$ and $W^{-}$ and are explained further in Sec. 6. Also shown are ALICE results at central rapidity and the output of the STARlight generator.
Figure 6: Differential cross-section multiplied by the branching fraction to dipions, in bins of rapidity, for $\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$ coherent photoproduction. The error bars represent the statistical uncertainty, while the bands correspond to the total uncertainty.
The results are given in Table 5 with the uncorrelated and correlated uncertainties quoted separately, and are shown in Figs. 5 and  6. For the $\rho$ meson, a comparison is made to ALICE data [25], to the predictions of Guzey, Kryshen and Zhalov (GKZ) [11], and to the results of the STARlight generator [37]. Further details of this calculation are provided in Sec. 6. The trend in the data for $\rho$ production is reproduced by the GKZ model, which takes account of elastic and inelastic nuclear shadowing. However, the GKZ model predicts a slightly higher overall cross-section. In other words, the data prefer more suppression than is given by the QCD effects considered in Ref. [11]. The measured cross-section roughly agrees with the STARlight generator although the $\rho$ measurement is systematically higher for rapidities above 3.5. However, no strong conclusions can be drawn since STARlight simply uses a nuclear form factor to scale the HERA photoproduction results (see the comment in Sec. 4 of Ref.[50]).
The results for the $\omega$ photoproduction cross-section have a strong fit-model dependence with the ratio $\sigma_{\rho}/\sigma_{\omega}$ varying between $9.8\pm 1.0$ (Eq. 4), $12.9\pm 1.3$ (Eq. 5) and $13.5\pm 1.4$ (Eq. 7). These results are in agreement with the determinations of this ratio by the ZEUS [51] and H1 [19] collaborations of $8.6\pm 1.0$ and $10.3^{+3.3}_{-2.3}$, respectively, using photoproduction on the proton at similar photon-proton centre-of-mass energies. The ZEUS collaboration measured the $\omega$ cross-section directly in the channel $\omega\rightarrow{{\pi}^{+}}{{\pi}^{-}}{{\pi}^{0}}$, while H1 measured it through its interference, as performed here, in the channel $\omega\rightarrow{{\pi}^{+}}{{\pi}^{-}}$.
@@ 6 Nuclear suppression factor
A comparison of the results on $\rho$ photoproduction on lead nuclei to photoproduction on the proton as measured by the H1 collaboration requires resolving the two-fold ambiguity in determining the photon emitter. However, at the highest rapidities, irrespective of nuclear suppression, the backward-going lead nucleus is almost always the photon emitter.
The differential cross-section for ${\rm PbPb}\rightarrow{\rm Pb}\rho{\rm Pb}$ can be expressed [11] in terms of the photoproduction cross-section $\gamma{\rm Pb}\rightarrow\rho{\rm Pb}$ using the photon flux $\mathrm{d}N_{\gamma}/\mathrm{d}k$ for the number of photons $N_{\gamma}$ of energy $k$:
@@
In the absence of nuclear effects, the $\rho$ photoproduction cross-section on the ion scales from the cross-section on the proton according to the number of nucleons and is given by the impulse approximation [52]. However, due to the large cross-section for $\rho$-nucleon scattering, as the meson traverses the nucleus, multiple interactions are possible. This can be taken into account in the Glauber model that considers elastic scatters, or extended to include inelastic scatters with a Gribov–Glauber model [50]. Details of these calculations are given in the Appendix. The predictions for the impulse approximation, Glauber and Gribov–Glauber models are shown in the left panel of Fig. 7, where it can be seen that the elastic and inelastic rescattering effects have a major impact on the result. The data are clearly inconsistent with the impulse approximation, while nuclear shadowing alone is also insufficient to provide a complete description. Including inelastic scattering improves the overall picture but some additional suppression is needed.
@@
The first term on the right-hand side of Eq. 10 only contributes 3% in the rapidity bin from 4.75 to 4.90, well within the systematic uncertainties. Ignoring this ‘+’ solution, Eq. 10 predicts a $\rho$ cross-section for PbPb collisions without nuclear suppression of $3330\pm 770$ mb, to be compared with the measured value of $625\pm 43\pm 44$ mb. In the perturbative regime where the Pomeron is equivalent to two gluons, the nuclear suppression factor has been defined [3] with respect to the impulse approximation using
@@
With the suppression factor measured relative to the impulse approximation, the fit returns values of $S_{\rm IA}^{-}=0.39\pm 0.03$ and $S_{\rm IA}^{+}=0.05\pm 0.11$ while $\lambda=0.6\pm 0.6$. This result shows that the events produced by higher-energy photons (lower-energy gluons) experience more nuclear suppression than those produced by lower-energy photons (higher energy gluons), as predicted by the Glauber and Gribov–Glauber calculations. A similar effect, albeit at a much reduced level in the pertubative ${J\mskip-3.0mu/\mskip-2.0mu\psi}$ process, has been reported in Refs. [3, 7] using ALICE and CMS UPC data [26, 31]. To take account of rescattering contributions, the nuclear suppression is calculated with respect to the Glauber approximation that includes elastic contributions predicted by QCD. This returns $S_{\rm Glauber}^{-}=0.78\pm 0.04$ and $S_{\rm Glauber}^{+}=0.30\pm 0.11$ while $\lambda=0.9\pm 0.9$. This shows that nonlinear effects are much more significant for the low-$x$ gluons. Finally, after including inelastic contributions using the Gribov–Glauber model, the relative suppression factors are $S_{\rm Gribov-Glauber}^{-}=0.84\pm 0.04$ and $S^{+}_{\rm Gribov-Glauber}=0.83\pm 0.06$ while $\lambda=0.7\pm 0.9$. The result is shown in the right panel of Fig. 7. The data is in closer agreement to the model although some additional suppression is still needed, which could be due to saturation effects or an increased dispersion of $P(\sigma)$.
@@ 7 Conclusions
The differential cross-section for the ${{\rho}^{0}}$ meson shows a slow rise with rapidity, in agreement with the GKZ model that includes linear and nonlinear QCD effects using a Gribov–Glauber calculation to describe multiple nucleon interactions. The observed cross-section is lower by about 30%, indicating that additional suppression is required. In PbPb interactions, there is a two-fold ambiguity in determining the photon emitter. However, by utilising the different cross-section dependencies with rapidity, there is sensitivity to both high and low $x$-values. Compared to the Gribov–Glauber model, suppression factors of $0.84\pm 0.04$ for values of $x\sim 5\times 10^{-3}$ and of $0.83\pm 0.06$ for $x\sim 5\times 10^{-6}$ are obtained. This additional suppression could be caused by saturation effects, but models that explicitly include saturation cannot be used to predict the cross-section since the ${{\rho}^{0}}$ meson is too light to apply perturbation theory. We thus rely on a phenomenological model that scales and extrapolates fits to fixed-target and H1 data for photoproduction on the proton. The LHCb photon-nucleon centre-of-mass energies in the range 170–760 GeV require an extrapolation of the H1 data, while energies in the range 5–23 GeV are only measured in fixed-target experiments, are less precise, and are also in a region where Reggeon exchange is important. Thus, new measurements of photoproduction on the proton are highly desirable at these energies and can be obtained at forward rapidities in proton-lead collisions at the LHC, or using fixed-target gaseous collisions at LHCb.
@@
The photon flux for quasi-real photons can be calculated in the equivalent photon approximation [53] as a function of $k$ and impact parameter $b$. Restricting $b$ to be greater than twice the radius of the ions, as is required in UPC, gives
@@
with $\zeta=2R_{\rm Pb}k/\upgamma$, where $R_{\rm Pb}=6.68\text{\,fm}$ [54] is the radius of the ion, $\upgamma$ is the nucleus Lorentz factor, $Z$ is the charge of the ion, $\alpha$ is the fine-structure constant and $K_{0},K_{1}$ are modified Bessel functions.
@@
The H1 collaboration measured $\rho$ production in the energy range $20<W<80\text{\,Ge\kern-1.00006ptV}$, while the range considered here is larger. For $W<20\text{\,Ge\kern-1.00006ptV}$, Reggeon as well as Pomeron contributions are important, while for $W>80\text{\,Ge\kern-1.00006ptV}$ an extrapolation of the HERA data is required. By fitting their data combined with fixed target data, the H1 collaboration found [19]
@@
Multiple interactions are taken into account using a Glauber calculation, assuming vector-meson dominance and using the optical theorem to relate the total cross-section to the imaginary part of the forward scattering amplitude. Following Ref. [50], the photoproduction cross-section on the nucleus can be written,
@@
The Glauber approximation only considers elastic scatters and needs to be extended to include inelastic scatters. This can be done with a Gribov–Glauber model as described in Ref. [50], and Eq. 22 is modified to
@@
The parameters, $C,\sigma_{0},\Omega$ are detailed in Ref. [50], where $C$ is a normalisation constant, the average total cross-section $\int\sigma P(\sigma)\mathrm{d}\sigma=\sigma_{\rho N}$, and the width of the distribution, $\Omega$ is constrained by data.
The uncertainties on the predictions are calculated as the quadratic sum of the uncertainty on the photon flux, obtained by changing the ion radius by 0.5 fm, and the uncertainty on $\mathrm{d}\sigma_{\gamma p\rightarrow p}/\mathrm{d}t(W,t=0)$, calculated from the uncertainties and correlations on the parameters of Eq. 21 given in Ref. [19]. For the Gribov–Glauber model, an additional uncertainty is added in quadrature, obtained by changing the dispersion of $P(\sigma)$ as described in Ref. [50].

### v2

@@ 4 Fit to the dipion spectrum
There is a long history of possible vector states in the high-mass region as discussed in the PDG review [PDG2024]. The $\rho^{\prime}$ resonance is measured to have a mass of $1350\pm 5\pm 20\text{\,Me\kern-1.00006ptV}$ and a width of $320\pm 10\pm 40\text{\,Me\kern-1.00006ptV}$, while the $\rho^{\prime\prime}$ resonance has a mass of $1790\pm 5\pm 20\text{\,Me\kern-1.00006ptV}$ and a width of $290\pm 10\pm 40\text{\,Me\kern-1.00006ptV}$, where the first uncertainty is the statistical precision of the fit result and the second is due to systematic effects, evaluated using the standard deviation of results obtained by fitting the model separately in each rapidity bin. The statistical significance of either or both resonances compared to the null hypothesis of Eq. 5 greatly exceeds the threshold for claiming an observation. However, it should be pointed out that improvements to the fit have been obtained with alternative parametrisations including using a form factor to multiply $\cal{BW}_{\rho}$, and including a contribution from a noninterfering $f_{2}(1270)$ meson.
These measurements are consistent with the values of the $\rho(1450)$ and $\rho(1700)$ mesons in the PDG [PDG2024], noting the role the choice of dynamic lineshapes can play. Further evidence for the existence of these particles was reported by the Babar collaboration using $e^{+}e^{-}\rightarrow\pi\pi(\gamma)$ data [BaBar:2012bdw] where the same qualitative features are observed as in the LHCb data. The production processes are similar with the main difference being that in photoproduction the photon is quasi-real while it is strongly virtual in $e^{+}e^{-}$ annihilation. The coupling between the photon and the pions is probed in both processes so the LHCb data may help clarify the resonant contributions, which have an impact in determining the hadronic corrections to the g2g-2 measurement [Davier:2019can].
@@ 5 Cross-section
[…]
@@
Table 5: Differential cross-section for the $\rho,\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$ mesons in bins of rapidity, using the model of Eq. 7. The first uncertainty includes statistical and uncorrelated systematics. The second uncertainty represents the correlated systematics.
@@
Figure 5: Differential cross-section for coherent $\rho$ photoproduction as a function of rapidity. The LHCb points use the model of Eq. 7 and the error bars represent the statistical and total uncertainties. The shaded band corresponds to the estimated uncertainty on the GKZ prediction. The contributions due to the photon emitted from each ion are labelled as $W^{+}$ and $W^{-}$ and are explained further in Sec. 6. Also shown are ALICE results at central rapidity and the output of the STARlight generator.
Figure 6: Differential cross-section multiplied by the branching fraction to dipions, in bins of rapidity, for $\omega,\rho^{\prime}$ and $\rho^{\prime\prime}$ coherent photoproduction, calculated using the model of Eq. 7. The error bars represent the statistical uncertainty, while the bands correspond to the total uncertainty.
The results are given in Table 5 with the uncorrelated and correlated uncertainties quoted separately, and are shown in Figs. 5 and  6. For the $\rho$ meson, a comparison is made to ALICE data [ALICE:2020ugp], to the predictions of Guzey, Kryshen and Zhalov (GKZ) [Guzey:2016piu], and to the results of the STARlight generator [Klein:2016yzr]. Further details of this calculation are provided in Sec. 6. The trend in the data for $\rho$ production is reproduced by the GKZ model, which takes account of elastic and inelastic nuclear shadowing. However, the GKZ model predicts a slightly higher overall cross-section. In other words, the data prefer more suppression than is given by the QCD effects considered in Ref. [Guzey:2016piu]. The measured cross-section roughly agrees with the STARlight generator although the $\rho$ measurement is systematically higher for rapidities above 3.5. However, no strong conclusions can be drawn since STARlight simply uses a nuclear form factor to scale the HERA photoproduction results (see the comment in Sec. 4 of Ref.[Frankfurt:2015cwa]).
The results for the $\omega$ photoproduction cross-section have a strong fit-model dependence with the ratio $\sigma_{\rho}/\sigma_{\omega}$ varying between $9.8\pm 1.0$ (Eq. 4), $12.9\pm 1.3$ (Eq. 5) and $13.5\pm 1.4$ (Eq. 7). These results are in agreement with the determinations of this ratio by the ZEUS [ZEUS:1996zse] and H1 [H1:2020lzc] collaborations of $8.6\pm 1.0$ and $10.3^{+3.3}_{-2.3}$, respectively, using photoproduction on the proton at similar photon-proton centre-of-mass energies. The ZEUS collaboration measured the $\omega$ cross-section directly in the channel $\omega\rightarrow{{\pi}^{+}}{{\pi}^{-}}{{\pi}^{0}}$, while H1 measured it through its interference, as performed here, in the channel $\omega\rightarrow{{\pi}^{+}}{{\pi}^{-}}$.
@@ 6 Nuclear suppression factor
Nuclear suppression effects are accessed by comparing the scaling between photoproduction on the proton, $\sigma_{\gamma p\rightarrow{{\rho}^{0}}p}$, and photoproduction on the nucleus, $\sigma_{\gamma{\rm Pb}\rightarrow{{\rho}^{0}}{\rm Pb}}$. First, the $\sigma_{\gamma p\rightarrow{{\rho}^{0}}p}$ results from the H1 [H1:2020lzc] experiment are scaled to $\sigma_{\gamma{\rm Pb}\rightarrow{{\rho}^{0}}{\rm Pb}}$ using three models described below: the impulse approximation, a Glauber model, and a Gribov–Glauber model. Then, after accounting for the photon fluxes from the lead nuclei, a prediction is obtained for the coherent production of $\rho$ mesons in PbPb collisions, which is compared to the experimental data.
The differential cross-section for ${\rm PbPb}\rightarrow{\rm Pb}\rho{\rm Pb}$ can be expressed [Guzey:2016piu] in terms of the photoproduction cross-section $\gamma{\rm Pb}\rightarrow\rho{\rm Pb}$ using the photon flux $\mathrm{d}N_{\gamma}/\mathrm{d}k$ for the number of photons $N_{\gamma}$ of energy $k$:
@@
In the absence of nuclear effects, the $\rho$ photoproduction cross-section on the ion scales from the cross-section on the proton according to the number of nucleons and is given by the impulse approximation [Chew:1952fca]. However, due to the large cross-section for $\rho$-nucleon scattering, as the meson traverses the nucleus, multiple interactions are possible. This can be taken into account in the Glauber model that considers elastic scatters, or extended to include inelastic scatters with a Gribov–Glauber model [Frankfurt:2015cwa]. Details of these calculations are given in the Appendix. The predictions for the impulse approximation, Glauber and Gribov–Glauber models are shown in the left panel of Fig. 7, where it can be seen that the elastic and inelastic rescattering effects have a major impact on the result. The data are clearly inconsistent with the impulse approximation, while nuclear shadowing alone is also insufficient to provide a complete description. Including inelastic scattering improves the overall picture but some additional suppression is needed.
@@
Equation 10 shows there is a two-fold ambiguity in determining the photon emitter. However, at the highest rapidities, irrespective of nuclear suppression, the backward-going lead nucleus is almost always the photon emitter. As a result, the first term on the right-hand side only contributes 3% in the rapidity bin from 4.75 to 4.90, well within the systematic uncertainties. Ignoring this ‘+’ solution, Eq. 10 predicts a $\rho$ cross-section for PbPb collisions without nuclear suppression of $3330\pm 770$ mb, to be compared with the measured value of $625\pm 43\pm 44$ mb. In the perturbative regime where the Pomeron is equivalent to two gluons, the nuclear suppression factor has been defined [Guzey:2020ntc] with respect to the impulse approximation using
@@
With the suppression factor measured relative to the impulse approximation, the fit returns values of $S_{\rm IA}^{-}=0.39\pm 0.03$ and $S_{\rm IA}^{+}=0.05\pm 0.11$ while $\lambda=0.6\pm 0.6$. This result shows that the events produced by higher-energy photons (lower-energy gluons) experience more nuclear suppression than those produced by lower-energy photons (higher energy gluons), as predicted by the Glauber and Gribov–Glauber calculations. A similar effect, albeit at a much reduced level in the pertubative $J\mskip-3.0mu/\mskip-2.0mu\psi$} process, has been reported in Refs. [Guzey:2020ntc, Mantysaari:2023xcu] using ALICE and CMS UPC data [ALICE:2023jgu, CMS:2023snh]. To take account of rescattering contributions, the nuclear suppression is calculated with respect to the Glauber approximation that includes elastic contributions predicted by QCD. This returns $S_{\rm Glauber}^{-}=0.78\pm 0.04$ and $S_{\rm Glauber}^{+}=0.30\pm 0.11$ while $\lambda=0.9\pm 0.9$. This shows that nonlinear effects are much more significant for the low-$x$ gluons. Finally, after including inelastic contributions using the Gribov–Glauber model, the relative suppression factors are $S_{\rm Gribov-Glauber}^{-}=0.84\pm 0.04$ and $S^{+}_{\rm Gribov-Glauber}=0.83\pm 0.06$ while $\lambda=0.7\pm 0.9$. The result is shown in the right panel of Fig. 7. The data is in closer agreement to the model although some additional suppression is still needed, which could be due to saturation effects or an increased dispersion of $P(\sigma)$.
@@ 7 Conclusions
The differential cross-section for the ${\rho}^{0}$} meson shows a slow rise with rapidity, in agreement with the GKZ model that includes linear and nonlinear QCD effects using a Gribov–Glauber calculation to describe multiple nucleon interactions. The observed cross-section is lower by about 30%, indicating that additional suppression is required. In PbPb interactions, there is a two-fold ambiguity in determining the photon emitter. However, by utilising the different cross-section dependencies with rapidity, there is sensitivity to both high and low $x$-values. Compared to the Gribov–Glauber model, suppression factors of $0.84\pm 0.04$ for values of $x\sim 5\times 10^{-3}$ and of $0.83\pm 0.06$ for $x\sim 5\times 10^{-6}$ are obtained. Additional suppression is observed and this could be caused by saturation effects. However, models that explicitly include saturation cannot be used to predict the cross-section since the ${\rho}^{0}$} meson is too light to apply perturbation theory. We thus rely on a phenomenological model that scales and extrapolates fits to fixed-target and H1 data for photoproduction on the proton. The LHCb photon-nucleon centre-of-mass energies in the range 170–760 GeV require an extrapolation of the H1 data, while energies in the range 5–23 GeV are only measured in fixed-target experiments, are less precise, and are also in a region where Reggeon exchange is important. Thus, new measurements of photoproduction on the proton are highly desirable at these energies and can be obtained at forward rapidities in proton-lead collisions at the LHC, or using fixed-target gaseous collisions at LHCb.
@@
The photon flux for quasi-real photons can be calculated in the equivalent photon approximation [Baur:2001jj] as a function of $k$ and impact parameter $b$. Restricting $b$ to be greater than twice the radius of the ions, as is required in UPC, gives
@@
with $\zeta=2R_{\rm Pb}k/\upgamma$, where $R_{\rm Pb}=6.68\text{\,fm}$ [Jones:2014aoa] is the radius of the ion, $\upgamma$ is the nucleus Lorentz factor, $Z$ is the charge of the ion, $\alpha$ is the fine-structure constant and $K_{0},K_{1}$ are modified Bessel functions.
@@
The H1 collaboration measured $\rho$ production in the energy range $20<W<80\text{\,Ge\kern-1.00006ptV}$, while the range considered here is larger. For $W<20\text{\,Ge\kern-1.00006ptV}$, Reggeon as well as Pomeron contributions are important, while for $W>80\text{\,Ge\kern-1.00006ptV}$ an extrapolation of the HERA data is required. By fitting their data combined with fixed target data, the H1 collaboration found [H1:2020lzc]
@@
Multiple interactions are taken into account using a Glauber calculation, assuming vector-meson dominance and using the optical theorem to relate the total cross-section to the imaginary part of the forward scattering amplitude. Following Ref. [Frankfurt:2015cwa], the photoproduction cross-section on the nucleus can be written,
@@
The Glauber approximation only considers elastic scatters and needs to be extended to include inelastic scatters. This can be done with a Gribov–Glauber model as described in Ref. [Frankfurt:2015cwa], and Eq. 22 is modified to
@@
The parameters, $C,\sigma_{0},\Omega$ are detailed in Ref. [Frankfurt:2015cwa], where $C$ is a normalisation constant, the average total cross-section $\int\sigma P(\sigma)\mathrm{d}\sigma=\sigma_{\rho N}$, and the width of the distribution, $\Omega$ is constrained by data.
The uncertainties on the predictions are calculated as the quadratic sum of the uncertainty on the photon flux, obtained by changing the ion radius by 0.5 fm, and the uncertainty on $\mathrm{d}\sigma_{\gamma p\rightarrow p}/\mathrm{d}t(W,t=0)$, calculated from the uncertainties and correlations on the parameters of Eq. 21 given in Ref. [H1:2020lzc]. For the Gribov–Glauber model, an additional uncertainty is added in quadrature, obtained by changing the dispersion of $P(\sigma)$ as described in Ref. [Frankfurt:2015cwa].